OpenStudy (callisto):

Plane geometry question #2 - centres in a triangle

OpenStudy (callisto):

I'm useless, sorry

OpenStudy (anonymous):

Orthoscenter is and intersection of the three altitudes of a triangle. To find it first find a line equation of each of the sides of the triangle and their respective perpendicular line pasing thru the respective vertice. Then just find their intersection point....

OpenStudy (callisto):

Good answer, but how can I find the equation the altitudes? I know one of them is x=18.. Sorry, I'm useless..

OpenStudy (anonymous):

 To find equation of for ex. up right side find the slope: (0-18)/(48-24)=-18/24=-3/4 so equation of line is (y-18) =-3/4(x-24) The perpendicular to this one passing thru the point (0,0): (y-0)=4/3(x-0) Do the same for the rest